OpenStudy (anonymous):

Simplify the following expression: 2x - 2y + 3z - 2x - y + 2z = ?

OpenStudy (anonymous):

do you have a calculator ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2x-2y+3z-2x-y+2z Just calculate it as you calculate normal expressions. The term with same letter can be calculate. 2x-2y+3z-2x-y+2z =(2x+2x)+(-2y-y)+(3z+2z) =4x-3y+5z

OpenStudy (anonymous):

@jadee27p

OpenStudy (anonymous):

can you help with another @chineseboy15

OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $6 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $78 for 1 adult ticket and 2 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 78 = one-halfx + (x - 6) What is the regular price of a child's ticket?

OpenStudy (texaschic101):

78 = 1/2x + x - 6 78 = 1/2x + 2/2x - 6 78 = 3/2x - 6 78 + 6 = 3/2x 84 = 3/2x 84/(3/2) = x 84 * 2/3 = x 168/3 = x 56 = x childs ticket is 6 less then adult....56 - 6 = 50 check... 1/2(56) + 1/2(2 * 50) = 78 28 + 1/2(100) = 78 28 + 50 = 78 78 = 78 (correct) regular adult tickets cost $56 regular childs tickets cost $50

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (anonymous):

Kimberley wants to plot a graph to show the price of different numbers of skateboards. She has to plot the following points: (1, 49.99), (2, 99.98), (3, 149.97), (4, 199.96) Which of the following best describes how the x-axis should be labeled? From 0 to 6 at intervals of 3 From 0 to 8 at intervals of 4 From 0 to 3 at intervals of 1 From 0 to 5 at intervals of 1 On this one i got D, am i right?

OpenStudy (texaschic101):

yes....I would label it 0 to 5 in intervals of 1

OpenStudy (anonymous):

can you help me with 2 more?

OpenStudy (texaschic101):

possibly...I can't make no promises

OpenStudy (anonymous):

The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank. 1 over 3π (x - 7)2 (x - 3) What does the factor (x - 7)2 represent? The square of the radius of the circular base of the tank The square of the height of the tank The area of the circular base of the tank The area of the curved sides of the tank

OpenStudy (texaschic101):

I am not the best at geometry...but I am guessing it is the square of the radius of the circular base of the tank.

OpenStudy (texaschic101):

I am pretty sure that is correct

OpenStudy (anonymous):

Matt needs to make $230 for a trip. He plans on selling 16 bags of snacks to raise money. The lowest acceptable price for him to sell these to make his goal is $______. This is the last one and for the one before this i got "The area of the circular base of the tank "

OpenStudy (texaschic101):

230/16 = 14.375 rounds to 14.38 check.. 16(14.38) >= 230 230.08 >= 230 (correct) The cheapest he can sell them is $14.38 in order to meet his goal as far as the one before....I think it is the square of the radius of the circular base. I think your answer is incorrect because the area of the circular base is pi * r^2

OpenStudy (anonymous):

for the first one you got 230 as the answer ?

OpenStudy (texaschic101):

for the first one, I got $14.38 as the answer

OpenStudy (texaschic101):

any questions ?